OpenStudy (tomboy213):

write an equation in point-slope form for the line through the given point with the given slope. (-4,6); m=3/4

OpenStudy (tomboy213):

Choices are: A. y-6=3/4x+4 B. y-6=3/4(x-4) C. y-6=3/4(x+4) D. y+6=3/4(x+4)

OpenStudy (mandywestside):

Try out Desmos! It's a great graphing calculator.

OpenStudy (tomboy213):

I think its c since the formula for point slop form is y-y1=m(x-x1)

Nnesha (nnesha):

^correct!

OpenStudy (tomboy213):

thank you